OpenStudy (anonymous):

The cylindrical walls of an open-top silo have an outside diameter of 5.0m, a thickness of 50cm, and a height of 10m. A can of paint covers 12m2, and cost $ 14.98. How much will it cost to paint the silo, including the inside, the outside, and the top of the walls?

OpenStudy (anonymous):

outside surface area=pi d*h=100pi inside surface area=pi*(4)*10=40pi so total area =140pi=439.822 549 dollars

OpenStudy (anonymous):

the real answer is $ 374.50 but i got $376.01

OpenStudy (anonymous):

sry calculation mistake outside are=50pi it you take base area=25/4pi so t.area=385/4pi 377.46$